OpenStudy (anonymous):

how do I simplify (27x^(1/2)y^(6))^(-1/3)

OpenStudy (jdoe0001):

http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (jdoe0001):

any part you need clearing out?

OpenStudy (anonymous):

my answer is 1/3x^(1/6) y^2

OpenStudy (anonymous):

but I need to show my steps and i don't know how

OpenStudy (anonymous):

hello?

OpenStudy (jdoe0001):

$$ \pmatrix{27x^{\frac{1}{2}}y^6}^{-\frac{1}{3}} $$

OpenStudy (jdoe0001):

start by the outtermost exponent, what would that give you?

OpenStudy (anonymous):

yup thats right

OpenStudy (anonymous):

-1/3?

OpenStudy (anonymous):

(27x^(1/2)y^(6))^(-1/3) can you rewrite it like shown on our paper use the equation botton

OpenStudy (jdoe0001):

yes, but what would say... \(a^{-1} = ? \)

OpenStudy (anonymous):

-a?

OpenStudy (anonymous):

ahh I am so confused

OpenStudy (jdoe0001):

ahemm, ok well have you checked http://www.math-play.com/image-exponents-rules.jpg yet?

OpenStudy (anonymous):

yeah

OpenStudy (jdoe0001):

because if you dunno the "exponent rules", then there's no use for this test

OpenStudy (anonymous):

ok so do I just apply each individual part to the rules?

OpenStudy (jdoe0001):

yes, starting from the outtermost exponent

OpenStudy (anonymous):

o

OpenStudy (jdoe0001):

so, from the sheet there in the link, look at the \(a^{-m} \)

OpenStudy (anonymous):

it'd be the -1/3

OpenStudy (jdoe0001):

in your case, yes

OpenStudy (jdoe0001):

so, what would that leave you with?

OpenStudy (anonymous):

3

OpenStudy (jdoe0001):

$$ a^{-\color{red}{2}} = \frac{1}{a^\color{red}{2}}\\ so\\ \pmatrix{27x^{\frac{1}{2}}y^6}^{-\color{red}{\frac{1}{3}}} \implies ? $$

OpenStudy (anonymous):

\[27x ^{1/2}y ^{6}/ (1/3)\]

OpenStudy (anonymous):

right?

OpenStudy (jdoe0001):

sorry I got a bit lagged

OpenStudy (jdoe0001):

well, I'll put this way $$ a^{-\color{red}{2}} = \frac{1}{a^\color{red}{2}}\\ so\\ \pmatrix{a}^{-\color{red}{\frac{1}{3}}} \implies ? $$

OpenStudy (anonymous):

\[1/(27x ^{1/2}y^6)^{-1/3}\]

OpenStudy (anonymous):

right?

OpenStudy (jdoe0001):

well, very close, yes, it becomes a fraction, but the "negative exponent" once at the denominator becomes "positive"

OpenStudy (anonymous):

oh whoops I forgot

OpenStudy (jdoe0001):

$$ a^{-\color{red}{2}} = \frac{1}{a^\color{red}{2}}\\ so\\ \pmatrix{27x^{\frac{1}{2}}y^6}^{-\color{red}{\frac{1}{3}}} \implies \cfrac{1}{\pmatrix{27x^{\frac{1}{2}}y^6}^{\color{red}{\frac{1}{3}}}} $$

OpenStudy (anonymous):

\[1/(27^{1/2}y ^{6})^{1/3}\]

OpenStudy (jdoe0001):

as far as having an exponent outside a parentheses, see the rules sheet that shows \(\large (a^m)^n = a^{mn}\)

OpenStudy (anonymous):

so it comes into parenthesis

OpenStudy (anonymous):

I would multiply it in but by which one?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

well -> \(\large (a^2b^3c)^4 = (a^{2\times4} b^{3\times 4} c^{1\times 4}) \)

OpenStudy (anonymous):

oh so all of them

OpenStudy (anonymous):

6*1/3 is 2 and 6* 1/2 is 3

OpenStudy (anonymous):

so my new equation would look like \[1/27x^3 y^2\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

oh wait I found an error

OpenStudy (jdoe0001):

ahem, you're multiplying by a fraction, \(\frac{1}{3}\) not by "3"

OpenStudy (anonymous):

it'd be 1/3 * 1/2 which =1/6 and 1/3*6=2

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

yeah so my new equation would be 1/27x^(1/6) y^2

OpenStudy (anonymous):

is that right?

OpenStudy (jdoe0001):

well, yes, just that, any number by itself, usually would have an exponent of 1, so 25 = 25^1 27 = 27^1 \(\large 27^{1\times \frac{1}{3}}\)

OpenStudy (anonymous):

ok now what?

OpenStudy (jdoe0001):

$$\huge { \pmatrix{27x^{\frac{1}{2}}y^6}^{-\color{red}{\frac{1}{3}}} \implies \cfrac{1}{\pmatrix{27x^{\frac{1}{2}}y^6}^{\color{red}{\frac{1}{3}}}}\\ \implies \cfrac{1}{27^{\color{red}{\frac{1}{3}}}x^{\color{red}{\frac{1}{6}}}y^\color{red}{2}} \implies \cfrac{1}{\sqrt[\color{red}{3}]{27}\sqrt[\color{red}{6}]{x}y^\color{red}{2}} } $$

OpenStudy (anonymous):

ok and then I get

OpenStudy (anonymous):

1/3x^(1/6)y^2

OpenStudy (anonymous):

right?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (jdoe0001):

if I can

OpenStudy (anonymous):

ok it looks like \[-5/\sqrt{5}-\sqrt{3}\]

OpenStudy (jdoe0001):

right, a fractional exponent, just means a root, so $$\large { a^{\frac{n}{m}} = \sqrt[m]{a^n} } $$

OpenStudy (jdoe0001):

is that => $$\large { \cfrac{-5}{\sqrt{5}-\sqrt{3}} }? $$